OpenStudy (anonymous):

One side of a square is 5n inches long. Another side is n + 16 inches long. Write an equation by setting these equal and solving for n. How long is a side of the square? One point for finding n and one point for finding the length of the side.

OpenStudy (anonymous):

5n = n + 16 4n =16 n=4 Length of S1 = 5n = 20 in S2 = n + 16 = 20 in